OpenStudy (anonymous):

What should be added to both sides of this equation to solve for the variable? s + (-10) = 14

OpenStudy (anonymous):

add 10 to both sides

OpenStudy (anonymous):

and then whats the answer

OpenStudy (anonymous):

s + (-10) = 14 , s - 10 = 14, s = 14 + 10

OpenStudy (anonymous):

what